OpenStudy (anonymous):

\[^{x^2}=49 solve the equation x^n=K\]

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

x^=49 solve the equation x^n= K

OpenStudy (anonymous):

\[x^2=49\implies x=7\text{ or } x =-7\]

OpenStudy (anonymous):

because you know that \(7^2=49\) and also \((-7)^2=49\)

OpenStudy (anonymous):

yes how did you that arrow thing

OpenStudy (anonymous):

latex magic

OpenStudy (anonymous):

right

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

if you ever want to see the code, right click on the equation and choose "show math as" then "latex"

OpenStudy (anonymous):

but to answer your question i typed \implies

OpenStudy (anonymous):

sorry needed some java, i dont know if i will be able to do all of this by monday, im stressing bad, graduation is Friday